OpenStudy (anonymous):

PRE-CALCULUS: Find the exact value of cos-1 the quantity square root of three divided by two.

OpenStudy (anonymous):

\[\cos^{-1} \left( \frac{ \sqrt{3} }{ 2} \right)\]

OpenStudy (anonymous):

we know that the range arccos(x) is 0 ≤ x ≤ π/2 ==> 0° ≤ x ≤ 180°. So we need to find the solution to cos(x) = -√3/2 this in the range of [0°, 180°]. If we construct a 30-60-90 triangle, we see that the 30° angle corresponds to this and so the angle required is x = 180° - 30° = 150°. Thus, arccos(-√3/2) = 150° = 5π/6. Answer provided by Yahoo Answers :)

OpenStudy (anonymous):

The answer amber provided was for \[\sf \large-\frac{\sqrt{3}}{2}\] by the way

OpenStudy (anonymous):

I think your answer is wrong, but the way you explain it is right.

OpenStudy (anonymous):

I know @LegendarySadist :)

OpenStudy (anonymous):

The easiest way to go about this sort of problem is simply to utilize the unit circle.

OpenStudy (anonymous):

For cos, the range is 0<y<pi, right?

OpenStudy (anonymous):

Yes, I actually based my answer on the unit circle. My answer is \[\frac{ \pi }{ 6 }\]

OpenStudy (campbell_st):

here is a triangle  so the cosine of which angle is \[\cos(?) = \frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

It is 30 degrees.

OpenStudy (anonymous):

right @campbell_st ?

OpenStudy (campbell_st):

that's correct...

OpenStudy (campbell_st):

you just need to remember the exact value triangles... I always think about the longest side is opposite the largest angle 2 = 90 1 = 30 shortest and smallest angle

OpenStudy (anonymous):

@campbell_st Oh, I don't that one. It is Geometry. Thanks for the help! :)